Electron plasma waves are propagated in a unIform plasma with KTe=100 ev, n=10^16m^-3 and B=0.if the frequency f is 1.1 G,What is the wavelength in cm?
1
Expert's answer
2020-06-17T09:29:16-0400

The average speed of electrons is


"v=\sqrt\frac{2K}{m}."

The wavelength is


"\lambda=\frac{v}{f}=\frac{1}{f}\sqrt\frac{2K}{m}=\\\space\\
=\frac{1}{1.1\cdot10^9}\sqrt\frac{2\cdot100\cdot1.62\cdot10^{-19}}{9.11\cdot10^{-31}}=5.39\cdot10^{-3}\text{ m}."